# Predictive maintenance solution for industrial systems - an unsupervised approach based on log periodic power law

### v2

@@ Abstract
A new unsupervised predictive maintenance analysis method based on the renormalization group approach used to discover critical behavior in complex systems has been proposed. The algorithm analyzes univariate time series and detects critical points based on a newly proposed theorem that identifies critical points using a Log Periodic Power Law function fits. Application of a new algorithm for predictive maintenance analysis of industrial data collected from reciprocating compressor systems is presented. Based on the knowledge of the dynamics of the analyzed compressor system, the proposed algorithm predicts valve and piston rod seal failures well in advance.
@@ I Introduction
Predicting failures, provided by PM can be very profitable for a company, under the condition that PM minimizes the number of false warnings (false positives) and maximizes the number of correctly predicted events (true positives). Creating a properly working PM process faces two main problems:
@@
    related to the definition of what is a failure in the considered technological process or in the IoT data,
@@
While the theoretical definition of failure is well known, in practice problems are encountered with its implementation. For various reasons (economic, productive, etc.) not every failure requires corrective action. Sometimes, a minor failure is not a good reason to stop a monitored machinery or a production process. This imprecise description of failure leads to substantial difficulties when attempting to use supervised methods to build a correct Predictive Maintenance process. This would suggest that unsupervised approaches may prove to be a better suited tool for building PM processes.
This article describes an unsupervised failure prediction method patent0 used to monitor reciprocating compressor systems based on a concept used, among others, in finance, called Log-Periodic Power Law (LPPL) proposed by the Johansen2000 and Sornette2003ab . However, due to the different nature of the data analyzed, the LPPL method cannot be applied in the same way as it was introduced for bubble (or anti-bubble) detection in economic time series. Due to the different definitions of PM failure in industrial applications, modifications need to be made to the LPPL method. In the case of data describing the financial market, the variable that directly characterizes changes in the system under analysis is examined. Such a variable is, for instance the index of the analyzed financial market, or directly the price of shares.
@@
In the description of the application of the presented method, which will be discussed in detail on the basis of the monitoring data from reciprocating compressors, such indirect data are, for instance, changes in the opening angle of suction or discharge valves inside the compression chamber as a function of the volume in the cylinder chamber expressed by the angle of rotation of the crankshaft. In other words, degradative (unmonitored) changes affect measured variable changes in a less visible (more distorted) ways than in the case of direct measurement of variables.
Every failure has a cause, an initiating event, let's call it an initial breakdown (IB). Using the IB concept, the failure analysis can be as follows:
  • until the IB occurs, the behavior of the machine is normal and shows no signs of failure.
  • After the IB occurs, noticeable changes begin to occur indicating future problems.
The idea of detecting the IB point in order to use it to predict failure is similar to the concept of detecting a point of trend change in a time series. Therefore, it is not necessary to predict the time of failure in the future. It is sufficient to determine whether it can be determined whether a given point in the time series is the initial (initiating) moment of IB or not. If it is known that the current point in the analyzed time series is an IB point, then, based on the knowledge of the dynamics of the monitored device, it is possible to identify the time window in the future (in units of operating time of the device) in which the failure will occur. It is impossible however, to predict the exact time of failure of the monitored device using the presented method. This is mainly due to unpredictable occurrences, such as: changes in the load of the device, periods of time when the device is switched off, etc.
The organization of the paper is as follows. Chapter II briefly describes the current status of unsupervised predictive maintenance in the industry. The chapter III answers the question of why the logarithmic-periodic oscillation detection method is suitable for detecting sudden failures in an industrial system. Chapter IV describes a numerical method for determining failure time points from the data. The part V introduces the data used for the analysis presented and the following section VI shows the results obtained using the new method. Finally, the part VII discusses the results, advantages and disadvantages of the proposed method.
@@ II Related works
Despite the appearance of clarity, the description of a method as unsupervised is sometimes used too hastily. As described in the section (I), it is not always possible to identify with sufficient precision the time of failure as well as its cause. Even after repairs, the process of determining the time of failure or its origin is fraught with difficulties. Therefore, before describing the status of work on unsupervised methods, it should be stated, that in this work the “unsupervised method” is understood as a solution that does not require the knowledge of labelled data of any kind as well as the lack of information concerning the input data's condition as either good/healthy or anomalous.
Following the work of survey_1 (which also includes references to other research works, as well as studies of specific solutions involving a broader understanding of the "unsupervised method"), the spectrum of available solutions can be generally divided into 2 categories.
@@
    Solutions involving the use of prediction techniques to predict points in the future and then compare them with actual data to detect anomalies, as shown in method_1 . Depending on the predictive solutions used, this category of methods requires a large amount of data. In this case, a separate problem is the accuracy of the prediction part, which is a supervised method. Hence the requirement to control it, which makes the whole solution complicated if one wants to use it in practical applications.
@@
    Solutions based on measurements of distance or similarity used to evaluate the degree of data anomaly. In this case, methods using clustering algorithms method_2 or determining similarity between time series or data are used. This approach sometimes leads to problems when encountering new data the like of which did not exist in the past. At times this renders them impossible to use for real-time data analysis.
The solution proposed in this paper is an attempt to solve the issues identified in both of the above categories: the problem of data demand and the problem associated with data behavior that is not present in the past.
@@
The purpose of the PM method is to provide predictions about future failures of the system described as a set of various components cooperating together. This section illustrates why the LPPL-based algorithm is applicable to failure prediction and describes the basics of the LPPL-based method.
@@
where the $\dot{h}\left(t\right)$ denotes the derivative of the function $h\left(t\right)$ in time $t$. The $\delta$ and $G$ are the parameters of model. The $G$ is a positive constant parameter. In the following, it is assumed that $h\left(t\right)$ corresponds to changes in the variable that is tracked and from which the failure is attempted to be predicted. The hazard function, based on conditional probability theory, measures the probability that the relevant variable will show signs of failure, given that the failure has not occurred before prior to time $t$.
@@
with $t_{c}$ as a constant corresponding to time in a future. As can be seen, the solutions found for $\delta=1$ (2) and for $\delta<1$ (3) do not converge for times $t>0$. Therefore, the time of failure cannot be determined from their behavior. The most interesting case is $\delta>1$ (4), where the solution has a converging point at a finite time $t_{c}$ in future. In our analysis, the time $t_{c}$ will denote a time of a potential failure and can be determined by the choice of $G=\left(\frac{1}{t_{c}}\right)^{\frac{1}{1-\delta}}$. Then $t_{c}=\frac{1}{\left(\delta-1\right)h\left(t_{0}\right)^{\delta-1}}$.
Let's assume that the degradation of a working part, whose deterioration (failure) is predicted, can be treated as a discontinuous stochastic process associated with a given monitored variable. To simplify the analysis, the second assumption is to treat the degradation changes of physical quantity $p\left(t\right)$ over time $t$ as a non-homogeneous Poisson process in which the changes occur according to the hazard rate function $h\left(t\right)$. The dynamics of such a process can be described by the equation
@@
$\log{\left[\frac{p\left(t\right)}{p\left(t\right)}\right]}=-\int_{t_{0}}^{t}h\left(u\right)du=P\left(t\right).$ (6)
@@
where $\eta=\frac{1}{1-\delta}$ and the $P\left(t\right)$ is shifted by the integration constant $\frac{h_{0}}{\eta+1}\left(t_{c}-t_{0}\right)^{\eta+1}$. The result obtained (7) coincides with work Ledoit2000 (compare with equation (3) in the reference 7)
@@
The basic assumption of the LPPL method is that, the described process is near the critical point of the second-order phase transition. In our case, this is the point in time at which the failure of the described system occurs. With this assumption, the final equation is obtained, the use of which for fitting is known in literature as the LPPL method Feigenbaum1996 , Sornette1996 .
Let $W\left(t\right)=\log\left(p\left(t\right)\right)$ and refer $p\left(t\right)$ to the variable by which our industrial system is analyzed as a function of time $t$. Let $t_{c}$ be the time of event that defines our phase transition in physical framework (critical point of time). Then the argument $x$ and the real function $F\left(x\right)$ are defined as
@@
The request of scaling invariance (9) with the general form of the solution postulated by eq. (11) can be rewritten as
@@
Therefore, the solution of equation (9) can be expressed as Nauenberg1975
@@
where $\Pi\left(\cdot\right)$ is a periodic function with period 11, i.e. $\Pi\left(y\right)=\Pi\left(y+1\right)$ . The index $n$ should be treated as one of the parameters characterizing the described physical system. Since $n$ and other parameters appearing in the function (16) are unknown, it is necessary to reformulate the function (16) in such a way that it can be used to fit existing data and thus determine whether a given point $t_{c}$ is a critical point.
An additional necessary condition that must be satisfied by function (16), or more precisely by its real part, is the trend that is determined by the power law ($n=0$), which is the leading order term and the oscillations associated with $n\neq 0$ will contribute as a next-to-leading order corrections.
@@
where, the variable $q$ denotes an parameter associated with the physical degradation mechanism of the described system.
@@
where the notation $c_{\pm 1}$ was used to denote ambiguity as to which coefficient is the second dominant one for k=1k=-1 or $k=+1$. Since only the real part of the expression is of interest (our measurements are real values), using the previous definition of the variable $F\left(x\right)$ (8) and generalizing our unknown parameters ($\mu$, $\nu$, $C$, $q$, $n$, $k$ and $W\left(t_{c}\right)$) by adding constant AA and phase $\Phi$ to the formula (24) to their new representations (AA, $B$, $m$, $C$, $\omega$, $\Phi$) one obtains the final formula which is referred to as a first-order model and used in LPPL literature Feigenbaum1996 ; Sornette1996
$\begin{split}W\left(t\right)\approx A+\left|t_{c}-t\right|^{m}\left[B+C_{1}\cos\left(\omega\log\left|t_{c}-t\right|+\Phi\right)\right].\end{split}$ (25)
@@
where $C_{1}=C\cos\left(\Phi\right)$ and $C_{2}=-C\sin\left(\Phi\right)$. Both equations (25, 26) can be used to find critical time points in the time series of input data that indicate component failures in the input data.
@@ IV Fitting method of the LPPL model to the data
Instead of trying to determine the critical time $t_{c}$ in the future, as is determined in the case of predictions of crashes in financial time series Sornette1996 ; Shu2019 ; Ledoit2000 , it is assumed that the failure happens "now". With this assumption, the calculations involved in fitting the function (26) to the data are performed for a range of time windows of different lengths from the past to "now". This corresponds to the hypothesis that the time point $t_{c}$ (time point of the phase transition) is "now" and corresponds to the last point in the input time series $t_{inp}$. Therefore, it is necessary to add to the set of parameters, an additional parameter specifying the length of the subset of time series points $t_{inp}$ preceding the time point $t_{c}$ for which the best fit of the function $W\left(t\right)$ (26) was found. The length of this subset is denoted as $l_{max}$.
In particular, our redefined set of arguments $x=t_{c}-t$ in the matching procedure is the set of points $\left<1,x_{l_{max}}\right$> in time units characteristic to the $t_{inp}$ series. The definition of the argument set $x=t_{c}-t$ in the matching procedure is replaced by the set of points $\left<1,x_{l_{max}}\right$>, where 11 corresponds to the present time and $l_{max}$ corresponds to the time of $l_{max}$ from the past. The index $l_{max}$ is the one of the parameters of the matching function (26).
As parameter fitting, the method described in the work of Shu2019 is used, appropriately modified for our purposes, i.e., by excluding the parameter corresponding to the critical time $t_{c}$ and adding the parameter $l_{max}$ to the fitting procedure.
@@
If both trends $T_{max}$ and $T_{min}$ determined for the function $f\left(x\right)$ have the same behavior (both increasing or decreasing) for the last point $\left(t_{n}-t_{n-1},y_{n-1}\right)$ of the series tsts, then the point $\left(t_{n},y_{n}\right)$ is the critical point for the series tsts. The trend of the series tsts will change to the opposite for next points $\left(t_{n+k},y_{n+k}\right)$ (where $k>0$) with respect to the trend for points preceding $t_{n}$.
In the presented method, each point satisfying fulfilling the requirements of Theorem (1) is treated as a point initiating a future failure (IB point). Determining the time of failure requires knowledge of the dynamics of the monitored system and will be discussed in the next section (VI).

### v3

@@ Abstract
A new unsupervised predictive maintenance analysis method based on the renormalization group approach used to discover critical behavior in complex systems has been proposed. The algorithm analyzes univariate time series and detects critical points based on a newly proposed theorem that identifies critical points using a Log Periodic Power Law function fit. Application of a new algorithm for predictive maintenance analysis of industrial data collected from reciprocating compressor systems is presented. Based on the knowledge of the dynamics of the analyzed compressor system, the proposed algorithm predicts valve and piston rod seal failures well in advance.
@@ I Introduction
Predicting failures, provided by PM, can be very profitable for a company, under the condition that PM minimizes the number of false warnings (false positives) and maximizes the number of correctly predicted events (true positives).
Let us define failure as the termination of the ability of a part of an industrial device to perform the required function with the required efficiency. Another aspect of a failure that needs to be specified is how it can manifest itself. A failure can develop gradually or appear as a sudden event. In this work we consider only those failures that develop over time.
Creating a properly working PM process faces two main problems:
@@
    related to the determination of time when a failure occurred in the considered technological process or IoT data,
@@
Proper identification of the failure in IoT data (data labeling process) is a very difficult task and usually requires very specialized knowledge. For various reasons (economic, technological), not every failure requires corrective action. This is because the required efficiency that defines a failure in the monitored unit may depend on the degree of reliance of the process on the considered failure. Sometimes, a minor failure is not a sufficient reason to stop a monitored machinery or a production process. The determination of so-called good IoT data, i.e., the data when the monitored device behaves normally, is also a data labeling process that requires expertise. In this work, every action that defines the state of the data is considered a data labeling process.
Attempts to build a solution based on supervised Machine Learning (ML) models will encounter the following problems:
  1. 1.
    the complexity of the data labeling process,
  2. 2.
    the model degradation: due to the evolution of IoT data, it is necessary to update the supervised model because the current one starts to behave incorrectly. Such an update forces the creation of new training data necessary for a new supervised model (new data labeling process),
  3. 3.
    the complex process of monitoring the predictions of the currently working model. This is a difficult task, especially when dealing with a small number of predicted failures monitoring0 .
In published works on failure prediction using supervised solutions review0 , the authors omit the need for a continuous data labeling process, as well as the need to monitor the degradation of the running model. Thus, an important economic part of PM projects based on supervised models is passed over in silence. In review articles review0 , methods based on Neural Network architecture are placed in the category of unsupervised methods. However, these methods also require labeled training data. In this article, the term “unsupervised method” is understood only as a solution that does not require any knowledge of labeled data and any information about the state of the input data as good/healthy or anomalous.
The complex and expensive cycle of building and implementing a supervised method for a predictive maintenance process, sketched above, strongly suggests that unsupervised approaches are a better tool for building PM processes. An unsupervised solution removes the necessity of:
  1. 1.
    a complex data labeling process running at all times as a part of a supervised model,
  2. 2.
    monitoring of the deterioration of the model quality and thus re-creating models by training them on updated training data.
This paper describes an unsupervised failure prediction method patent0 used to monitor reciprocating compressor systems based on a concept called Log-Periodic Power Law (LPPL) proposed in Johansen2000 and Sornette2003ab . The LPPL mechanism is used to predict disasters in financial data finance0 , earthquakes earthquakes0 , avalanches and similar phenomena avalanches0 .
However, due to the different nature of the data analyzed, the LPPL method cannot be applied in the same way as for bubble (or anti-bubble) detection in economic time series. Due to the different definitions of PM failure in industrial applications, modifications need to be made to the LPPL method. In a case of data describing the financial market, the variable that directly characterizes changes in the system under analysis is examined. Such a variable is, for example, the index of the analyzed financial market or the prices of shares.
@@
In the description of the application of the presented method, such indirect data are, for instance, changes in the opening angle of suction or discharge valves inside the compression chamber as a function of the volume in the cylinder chamber expressed by the angle of rotation of the crankshaft. In other words, degradative (unmonitored) changes affect measured variable changes in a less visible (more distorted) ways than in case of a direct measurement of variables.
The principle of operation of the proposed method is based on the detection of such behavior of the analyzed data, which is in accordance with the logarithmic periodic power law. Finding such a functional behavior in the data is equivalent to finding the critical point of the system. It corresponds to the situation in which the system begins to work with decreasing efficiency, which is a signature of future failure. The point in the time series thus identified defines the start of the countdown to the predicted failure. Let’s call the trend change point the initial breakdown (IB). The concept of detecting an IB point for use in predicting failure is similar to the concept of detecting a trend change point in a time series. Therefore, it is not necessary to predict the time of failure in the future. All that is needed is to determine whether a given point in the time series is or is not an IB starting point. If it is known that the current point in the analyzed time series is an IB point, then based on the knowledge of the dynamics of the monitored device, it is possible to determine the time window in the future (in units of the device’s operating time) in which the failure will occur.
It should be emphasized that the determination of the IB point is not equivalent to the detection of anomalies. Anomalies in the behavior of the unit should be expected only in the time window of the expected failure. The concept of a time window in the future during which a failure is predicted is due to the fact that it is not possible to predict the exact time of failure of the monitored device using the presented method. This is mainly due to unpredictable events, such as: changes in the load of the device, periods when the device is off or independent maintenance activities.
The organization of the paper is as follows. Chapter II briefly describes the current status of unsupervised predictive maintenance in the industry. Chapter III answers the question of why the logarithmic-periodic oscillation detection method is suitable for detecting sudden failures in an industrial system. Chapter IV describes a numerical method for determining failure time points from the data. The part V introduces the data used for the analysis presented and the following chapter VI shows the results obtained using the new method and compares these results with those obtained using online statistical method. Finally, the part VII discusses the results and summarizes the advantages and disadvantages of the proposed method. This section also briefly outlines the criterion for selecting data to be used by this method.
@@ II Related works
Given the introduced definition of “unsupervised method”, as a methodology that does not require knowledge of the state of the input data, the list of existing methods for failure prediction is very limited. Following the work of survey_1 (which also includes references to other research works, as well as studies of specific solutions involving a broader understanding of the "unsupervised method"), the spectrum of available solutions can be generally divided into 2 categories.
@@
    Solutions involving the use of prediction techniques to predict points in the future and then compare them with actual data to detect anomalies, as shown in method_1 . Depending on predictive solutions used, this category of methods requires a large amount of data. In this case, a separate problem is the accuracy of the prediction part. Where the prediction part is a supervised method. Hence the requirement to control it, which makes the whole solution complicated if one wants to use it in practical applications.
@@
    Solutions based on measurements of distance or similarity used to evaluate the degree of data anomaly. In this case, methods using clustering algorithms method_2 or determining similarity between time series or data are used. This approach sometimes leads to problems when encountering new data that did not exist in the past. It happens that it makes it impossible to use them for real-time data analysis.
The solution proposed in this paper is an attempt to solve the issues identified in both of the above categories: the problem of demand for large amounts of data and the problem of data changes that do not occured in the past.
The essence of the proposed method is to determine the initial breakdown (IB) points in the time series of input data. Thus, it is natural to compare the proposed model with online methods that determine trend change points in time series in fully unspervised mode. Latest advances in statistical online trend change point detection can be found in studies changepoint1 ; changepoint2 . The difference between the statistical methods of determining the trend change points and the IB points determined by the proposed method is that in the proposed solution we are based on the search for a specific functional behavior (pattern) in the data. As will be shown later, this pattern is characteristic for physical processes leading to a degradation of the monitored components of the device A comparison of the results of the two methods, based on the same input data (statistical changepoint2 ; changepoint3 and the proposed model), will be presented and discussed in the section VI.
@@
The purpose of the PM method is to provide predictions for future failures of the system described as a set of various components cooperating together. This section illustrates why the LPPL-based algorithm is applicable to failure prediction and describes the basics of the LPPL-based method.
@@
where the $\dot{h}\left(t\right)$ denotes the derivative of the function $h\left(t\right)$ in time $t$. The $\delta$ and $G$ are the parameters of model. The $G$ is a positive constant parameter. In the following, it is assumed that $h\left(t\right)$ corresponds to changes in the variable that is tracked and from which the failure is attempted to be predicted. The hazard function, based on conditional probability theory, measures the probability that the relevant variable will show signs of a critical behaviour, given that the critical behaviour has not occurred prior to time $t$.
@@
with $t_{c}$ as a constant corresponding to time in a future. As can be seen, the solutions found for $\delta=1$ (2) and for $\delta<1$ (3) do not converge for times $t>0$. Therefore, the time of critical behaviour cannot be determined. The most interesting case is $\delta>1$ (4), where the solution has a converging point at a finite time $t_{c}$ in future. In our analysis, the time $t_{c}$ will denote a critical time point and can be determined by a choice of $G=\left(\frac{1}{t_{c}}\right)^{\frac{1}{1-\delta}}$. Then $t_{c}=\frac{1}{\left(\delta-1\right)h\left(t_{0}\right)^{\delta-1}}$.
Let us assume that the degradation of a working part, can be treated as a discontinuous stochastic process associated with a given monitored variable. To simplify the analysis, the second assumption is to treat the degradation changes of physical quantity $p\left(t\right)$ over time $t$ as a non-homogeneous Poisson process in which the changes occur according to the hazard rate function $h\left(t\right)$. The dynamics of such a process can be described by the equation
@@
$\log{\left[\frac{p\left(t\right)}{p\left(t_{0}\right)}\right]}=-\int_{t_{0}}^{t}h\left(u\right)du=P\left(t\right).$ (6)
@@
where $\eta=\frac{1}{1-\delta}$ and the $P\left(t\right)$ is shifted by the integration constant $\frac{h_{0}}{\eta+1}\left(t_{c}-t_{0}\right)^{\eta+1}$. The result obtained (7) coincides with work Ledoit2000 (compare with equation (3) in the reference Ledoit2000 )
@@
The basic assumption of the LPPL method is that, the described process is near the critical point of the second-order phase transition. With this assumption, the final equation is obtained, the use of which for fitting is known in literature as the LPPL method Feigenbaum1996 ; Sornette1996 .
Let $W\left(t\right)=\log\left(p\left(t\right)\right)$ and $p\left(t\right)$ corresponds to the variable by which our industrial system is analyzed as a function of time $t$. Let $t_{c}$ be the time of event that defines our phase transition in physical framework (critical time point). Then the argument $x$ and the real function $F\left(x\right)$ are defined as
@@
The condition of scaling invariance (9) with the general form of the solution postulated by eq. (11) can be rewritten as
@@
Therefore, the solution of equation (9) can be expressed, according to Nauenberg1975 , as
@@
where $\Pi\left(\cdot\right)$ is a periodic function with period $1$, i.e. $\Pi\left(y\right)=\Pi\left(y+1\right)$ . The index $n$ should be treated as one of the parameters characterizing the described physical system. Since $n$ and other parameters appearing in the function (16) are unknown, it is necessary to reformulate the function (16) in such a way that it can be used to fit existing data and thus determine whether a given point $t_{c}$ is a critical point.
An additional necessary condition that must be satisfied by function (16), or more precisely by its real part, is the trend that is determined by the power law ($n=0$), which is the leading order term and the oscillations associated with $n\neq 0$ will contribute as next-to-leading order corrections.
@@
where, the variable $q$ denotes a parameter associated with the physical degradation mechanism of the described system.
@@
where the notation $c_{\pm 1}$ was used to denote ambiguity as to which coefficient is the second dominant one for k=1k=-1 or $k=+1$. Since only the real part of the expression is of interest (our measurements are real values), using the previous definition of the variable $F\left(x\right)$ (8) and generalizing our unknown parameters ($\mu$, $\nu$, $C$, $q$, $n$, $k$ and $W\left(t_{c}\right)$) by adding constant $A$ and phase $\Phi$ to the formula (24) to their new representations ($A$, $B$, $m$, $C$, $\omega$, $\Phi$) one obtains the final formula which is referred to as a first-order model and used in LPPL literature Feigenbaum1996 ; Sornette1996
$\begin{split}W\left(t\right)\approx A+\left|t_{c}-t\right|^{m}\left[B+C\cos\left(\omega\log\left|t_{c}-t\right|+\Phi\right)\right].\end{split}$ (25)
@@
where $C_{1}=C\cos\left(\Phi\right)$ and $C_{2}=-C\sin\left(\Phi\right)$. Both equations (25, 26) can be used to find critical time points in the time series of input data.
@@ IV Fitting method of the LPPL model to the data
Instead of trying to determine the critical time $t_{c}$ in the future, as is determined in the case of predictions of crashes in financial time series Sornette1996 ; Shu2019 ; Ledoit2000 , it is assumed that the critical time point is "now". With this assumption, the calculations involved in fitting the function (26) to the data are performed for a range of time windows of different lengths from the past to "now". This corresponds to the hypothesis that the time point $t_{c}$ (time point of the phase transition) is "now" and corresponds to the last point in the input time series $t_{inp}$. Therefore, it is necessary to add to the set of parameters, an additional parameter specifying the length of the subset of time series points $t_{inp}$ preceding the time point $t_{c}$ for which the best fit of the function $W\left(t\right)$ (26) was found. The length of this subset is denoted as $l_{max}$.
In particular, our redefined set of arguments $x=t_{c}-t$ in the matching procedure is the set of points $\left<1,x_{l_{max}}\right$> in time units characteristic to the $t_{inp}$ series. The definition of the argument set $x=t_{c}-t$ in the matching procedure is replaced by the set of points $\left<1,x_{l_{max}}\right$>, where $1$ corresponds to the present time and $l_{max}$ corresponds to the time of $l_{max}$ from the past. The index $l_{max}$ is the one of the parameters of the matching function (26).
As parameter fitting, the method described in the work of Shu2019 is used, appropriately modified for our purposes, i.e., by excluding the parameter corresponding to the $t_{c}$ and adding the parameter $l_{max}$ to the fitting procedure.
@@
If both trends $T_{max}$ and $T_{min}$ determined for the function $f\left(x\right)$ have the same behavior (both increasing or decreasing) for the last point $\left(t_{n}-t_{n-1},y_{n-1}\right)$ of the series $ts$, then the point $\left(t_{n},y_{n}\right)$ is the critical point for the series $ts$. The trend of the series $ts$ will change to the opposite for next points $\left(t_{n+k},y_{n+k}\right)$ (where $k>0$) with respect to the trend for points preceding $t_{n}$.
In the presented method, any point that satisfies the Theorem (1) is treated as a point initiating a future failure - the initial breakdown (IB) point.
The formalism described above can be summarized as follows: the appearance of a failure due to deterioration of components in a complex system is preceded by the appearance of a phase transition of the 2nd kind, defined by point IB (critical time point). From this time point, the process of failure development begins. This period (after the time defined by the IB point) can be monitored by detecting anomalies. Anomalies (as a result of reduced process efficiency) occur when the effects of the failure become large enough and begin to affect the behavior of the monitored device.
Determining the time of failure requires knowledge of the dynamics of the monitored system and will be discussed in section VI.